If the area of a room is 55 square feet, find the length and width of the room.
Can u show me how i would solve this please?
THANK YOU FOR YOUR TIME.

Answer by checkley71(8403)

AREA OF THE ROOM = LENGTH TIMES WIDTH OR L*W=55 THUS THE ROOM CAN BE 11*5=55 OR 10*5.5=55 OR 7.4161984*7.426984 OR 8*6.875 OR 9*6.1111111 OR MANY OTHER COMBINATIONS THAT WHEN MULTIPLIED TOGETHER = 55